OpenStudy (anonymous):

find the values of the variables: [15 8] [2 5] [13 3] - = [4x 0] [x 3y+6] [12 4y+8]

OpenStudy (unklerhaukus):

matrix subtraction?

OpenStudy (unklerhaukus):

\[\begin{bmatrix}15,&8\\4x&0\end{bmatrix}-\begin{bmatrix}2,&5\\x,&3y+6\end{bmatrix} =\begin{bmatrix}15-2,&8-5\\4x-x,&0-(3y+6)\end{bmatrix}=\begin{bmatrix}13,&3\\12,&4y+8\end{bmatrix}\]

OpenStudy (unklerhaukus):

If you look at the bottom-left element you'll see that \(4x-x = 12\) this can be solved for \(x\). And if you look at the bottom right element you'll see \(0-(3y+6)=4y+8\) this can be solved for \(y\).